OpenStudy (anonymous):

solve the following equations: 1. 5^(x+1)=25^(1-x) 2. log3(x-1)=log3(2x/(1-x))

OpenStudy (anonymous):

for # 1 i got x=1/2 but im not confident

OpenStudy (anonymous):

5^(x+1)=5^(2(1-x)) x+1=2-2x 3x=1 x=1/2

OpenStudy (anonymous):

2) (x-1)=2x/(1-x) (1-x)^2+2x=0 1+x^2=0 so x is imaginory...no real solution..

OpenStudy (anonymous):

damn you are fast. its the same as using log right?

OpenStudy (anonymous):

of which one are u talking??

OpenStudy (anonymous):

the first and the second one the log cancle? right?

OpenStudy (anonymous):

oh isnt it suppose to be (1-x)^2-2X=0